# A Bayesian Nonparametric Stochastic Block Model for Directed Acyclic Graphs

### v2

@@ Abstract
Random graphs have been widely used in statistics, for example in network and social interaction analysis. In some applications, data may contain an inherent hierarchical ordering among its vertices, which prevents any directed edge between pairs of vertices that do not respect this order. For example, in bibliometrics, older papers cannot cite newer ones. In such situations, the resulting graph forms a Directed Acyclic Graph. In this article, we propose an extension of the popular Stochastic Block Model (SBM) to account for the presence of a latent hierarchical ordering in the data. The proposed approach includes a topological ordering in the likelihood of the model, which allows a directed edge to have positive probability only if the corresponding pair of vertices respect the order. This latent ordering is treated as an unknown parameter and endowed with a prior distribution. We describe how to formalize the model and perform posterior inference for a Bayesian nonparametric version of the SBM in which both the hierarchical ordering and the number of latent blocks are learnt from the data. Finally, an illustration with a real-world dataset from bibliometrics is presented. Additional supplementary materials are available online.
@@ 1 Introduction
Random graphs have been widely used in statistics and computer science over the last years to model network, interaction and bibliometric data, just to mention a few among many fields of applications. For example, in social networks data analysis, datasets collecting friendships or exchange of messages among users can be modelled as random graphs, having users as vertices, and edges between two users if they are friends or have exchanged messages. As another example, citation datasets in biliometrics analysis can be modelled as random graphs, having articles as vertices, and directed edges between a pair of articles if the former article cites the latter one.
@@
In some applications, the dataset, and corresponding graph, can contain an inherent hierarchical order among its vertices. This means that directed edges can be present only between vertices in which the former appears before the latter one in this order. As an example, when modelling citation networks among journal articles, (Ji and Jin, 2016), articles published many years ago cannot cite recent articles. As another example, in a network having the employees of a big company as vertices and directed edges from the supervisor to the supervisee, a directed path from one vertex to another can be present only if the former is higher than the latter one in the company hierarchy. Similarly, the same structure appears in a dataset having directed edges from PhD supervisors to their student. In all these applications, the observed graph is a directed graph that cannot display any directed cycles (i.e. any directed path in the graph having the same initial and final vertex). Hence, it is a Directed Acyclic Graph (DAG) and includes an inherent unknown topological ordering among its vertices, i.e. a linear ordering $\prec$ of the vertex set $\mathcal{V}$ such that every pair $(p,q)\in\mathcal{V}\times\mathcal{V}$ can belong to the edge set $E$ only if $p\prec q$.
@@
When analyzing a given network dataset that is a DAG, if the latent topological ordering is not properly accounted for, the statistical models can assign positive probability to ‘impossible’ edges, i.e. directed edges in which the first vertex appears later than the second one in the topological order. If this is the case for many edges, then the resulting estimates from the model can deteriorate dramatically. This is especially true for Bayesian models, and the same phenomenon arises other statistical contexts, for example in presence of structural zeros when modeling tabular data, Manrique-Vallier and Reiter (2014), in which assigning, whatever small, positive probability to ‘impossible’ cells can deteriorate estimates and bias results.
@@
In this work, we propose an approach to extend Bayesian models for random graphs to graphs that are DAGs and contain a latent topological ordering among their vertices. Specifically, we focus on extending the popular Stochastic Block Model (SBM), (Holland et al., 1983), to the case of an observed DAG. The SBM is among the most popular random graphs models for vertex clustering and community detection. This model was initially introduced in (Holland et al., 1983) and (Wang and Wong, 1987) and its statistical formalization as latent models was then described in (Snijders and Nowicki, 1997) and (Nowicki and Snijders, 2001). The model further gained popularity with the incorporation of degree correction (Karrer and Newman, 2011). In online Appendices A.1 and A.2, we provide a brief review of this model, its generalizations and approaches to posterior inference. For recent comprehensive reviews, the reader is also referred to (Abbe, 2018) and (Lee and Wilkinson, 2019), which focus on theoretical results and modeling approaches, respectively. When extending the SBM to model a given DAG, the underlying topological order is unknown, and not necessarily unique. Therefore, it will be included in the likelihood of the model as one of the unknown parameters, endowed with a prior and inferred a posteriori, together with the community memberships and other parameters of the model.
@@
The background literature regarding the use of the SBM for DAGs is quite limited and mainly consists of an unpublished technical report by Lee and Wilkinson (2018), which only presents a preliminary data analysis of a DAG graph using a SBM. Not only was degree correction not incorporated, but there was also no formal consideration of the issue of the number of blocks. Instead, the model was merely fitted with different numbers of blocks separately with no model selection, thus not justifying the interpretations of their results.
@@
The issue of the number of communities or blocks is common with SBMs and clustering in general, and has been reviewed in Table 1 of Lee and Wilkinson (2019). One common approach to this issue is the use of Bayesian nonparametrics, which assigns a prior for the allocation vector and allows the posterior sampler to infer the number of blocks from the data. Kemp et al. (2006) and Schmidt and Mørup (2013) initially considered the infinite relational model, which is a variant of the plain SBM, using a Dirichlet process prior, while Geng et al. (2019) extends this model by the use of a finite regime Pitman-Yor (PY) process (Pitman and Yor, 1997), which is reviewed in the online Appendix A.3. A further generalization is in Legramanti et al. (2022), in which the Gibbs-type prior formulation is used instead of the Pitman-Yor prior, for an application to criminal networks.
@@
In the proposed model, we also use a Bayesian nonparametric version of the SBM, but by adopting the PY process in a different way. This process allows two possible regimes: one in which the total number of blocks is finite (and unknown) and another one in which it increases the sample size. While other works such as Geng et al. (2019) have used similar Bayesian nonparametric formulation for the block membership, they usually focused only on the first of the two regimes. These two regimes imply different asymptotic behaviours of the number of blocks, and the model might be misspecified if only one regime is assumed before fitting to the data. To circumvent this, the choice of regime becomes another part of the proposed model, with a model selection step based on Gibbs variable selection (Carlin and Chib, 1995) embedded in the MCMC sampler.
@@
The proposed approach is illustrated both using simulated data and real datasets from bibliometrics. Specifically, the proposed model including the topological order is compared to a SBM for directed graphs, which does not account for the latent order in the likelihood. The simulation study shows that the latter model assigns positive probability to many ‘impossible’ edges, and brings about substantial biases in the estimates of the overall number of blocks. Then, we present an illustration with two datasets of citation networks, in which it is shown how the model can infer the community structure and latent hierarchical order among vertices.
@@
The rest of this article is organized as follows. The SBM for DAGs is presented in Section 2. Section 3 presents both simulated, Section 3.1, and real data, Section 3.2, illustrations. A final discussion Section 4 concludes the article. The online Appendices contain reviews of the SBM and PY process, a detailed description of the MCMC sampler to perform posterior inference for the SBM for DAG, including both a collapsed and an uncollapsed version of the algorithm, and detailed descriptions of the split and merge proposal and model selection step, other than including some additional information on the likelihood derivation and plots from the real-data illustration.
@@ 2 Model
To introduce the model and some notation that will be used in the rest of the paper, let us consider a directed network to be represented as a directed multi-graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the vertex set and $\mathcal{E}$ is the edge set. The size of $\mathcal{V}$, denoted by $n:=|\mathcal{V}|$, is the number of vertices of the graph $\mathcal{G}$. The $n\times{}n$ adjacency matrix of the graph is denoted by $\mathbf{Y}=(Y_{pq})_{1\leq p,q\leq n}$. If there are $y$ directed edges from vertex $p$ to vertex $q$ i.e. $y$ copies of $(p,q)$ in $\mathcal{E}$, then $Y_{pq}=y$, otherwise $Y_{pq}=0$. We also define $\mathbf{Z}_{n}=(Z_{1},\ldots,Z_{n})$ to be the allocation vector of length $n$, where $Z_{p}$ is a label associated to vertex $p$. Two vertices, $p$ and $q$, belong to the same group if and only if $Z_{p}=Z_{q}$. Essentially, $\mathbf{Z}_{n}$ represents the group memberships of the vertices. We assume there are $K_{n}>1$ unique labels displayed in $\mathbf{Z}_{n}$, denoted $(Z_{1}^{*},\ldots,Z_{K_{n}}^{*})$ hence $K_{n}$ groups of vertices. For notation easiness, now on, we will often write $\mathbb{I}(Z_{p}=i)$, in place of $\mathbb{I}(Z_{p}=Z_{i}^{*})$ and, similarly, given a matrix $\mathbf{C}$ and the event $Z_{p}=Z_{k}^{*}$, $C_{Z_{p}j}$ will denote element $C_{kj}$. Here and in the rest of the paper, $\mathbb{I}(A)$ denotes the indicator function of the event $A$, i.e. $\mathbb{I}(A)=1$ if $A$ is true, 00 otherwise.
@@
The main idea of the classic SBM is that the number of edges $Y_{pq}$ from vertex $p$ to vertex $q$ is independent of that of any other dyad, conditional on their group memberships, $Z_{p}$ and $Z_{q}$, (Holland et al., 1983, Nowicki and Snijders, 2001). Moreover, we will consider the degree-corrected version of the SBM, (Karrer and Newman, 2011), which takes into account the degree heterogeneity of the vertices within the same group, by introducing the vector $\boldsymbol{\xi}=(\xi_{1},\ldots,\xi_{n})$ of vertex specific parameters, where $\xi_{p}>0$ is the degree correction factor for vertex $p$. Mathematically, the number of edges between vertex $p$ and $q$ is modeled as
$Y_{pq}|\mathbf{C},\mathbf{Z}_{n},\boldsymbol{\xi}~{}\sim~{}\text{Poisson}(\xi_{p}\xi_{q}C_{Z_{p}Z_{q}}),$ (1)
where $\mathbf{C}=(C_{ij})\in\mathbb{R}_{+}^{K_{n}\times K_{n}}$ is the block matrix. The choice of the Poisson distribution, instead of the Bernoulli distribution, is due to a few reasons, namely, the natural extension to multi-graphs where $\mathbf{Y}$ can take non-negative integer values, the asymptotic equivalence between the edge probability and the expected number of edges for large sparse graphs (Karrer and Newman, 2011), and also the computational simplicity that will be illustrated in our model, especially when considering the degree-corrected version of the model.
@@
In the rest of this work, we will consider a graph $\mathcal{G}$ that is a DAG, and extend the classic SBM to model a DAG. As the graph is directed, $Y_{pq}$ is not necessarily the same as $Y_{qp}$. However, due to the acyclic nature of $\mathcal{G}$, the combinations of the possible values of $(Y_{pq},Y_{qp})$ are restricted to be $(0,0)$, $(0,y)$ or $(y,0)$, for some positive integer $y\in\mathbb{N}$. Also, $Y_{pp}=0$ for all $p=1,2,\ldots,n$, or equivalently the major diagonal contains all zeros. To utilize a unique feature of DAGs, as discussed in Section 1, we define $\boldsymbol{\sigma}=(\sigma_{1},\ldots,\sigma_{n})$ as the $n$-vector random variable that represents the ordering of $\mathcal{G}$, with the collection of all permutations of $\{1,2,\ldots,n\}$ as the sample space. A value of $\boldsymbol{\sigma}$ is deemed to be topological with respect to $\mathcal{G}$ if it satisfies the definition of topological ordering, i.e. if vertex $p$ topologically precedes vertex $q$, then there cannot be edges from vertex $q$ to vertex $p$. We define several quantities implied from $\boldsymbol{\sigma}$. First, $\boldsymbol{\phi}=(\phi_{1},\ldots,\phi_{n})$ is the “inverse” of $\boldsymbol{\sigma}$, which means that if vertex $p$ is the $r$-th vertex in the topological ordering, we have $\sigma_{r}=p$ and $\phi_{p}=r$. Essentially, $\boldsymbol{\phi}$ contains the position of each vertex in $\boldsymbol{\sigma}$. For convenience, without confusion, we write $\boldsymbol{\sigma}^{-1}=\boldsymbol{\phi}$ and $\boldsymbol{\phi}^{-1}=\boldsymbol{\sigma}$. Second, $\mathbf{Z}_{n}^{\boldsymbol{\sigma}}=\left(Z^{\boldsymbol{\sigma}}_{1},Z^{\boldsymbol{\sigma}}_{2},\ldots,Z^{\boldsymbol{\sigma}}_{n}\right)$ is the reordered allocation vector, where $Z^{\boldsymbol{\sigma}}_{p}=Z_{\sigma_{p}}$, where $Z_{\sigma_{p}}$ comes from the allocation vector $\mathbf{Z}_{n}$. In the same way, if we define $\boldsymbol{\xi}=(\xi_{1},\ldots,\xi_{n})$ as the vector of degree correction parameters, $\boldsymbol{\xi}^{\boldsymbol{\sigma}}=\left(\xi^{\boldsymbol{\sigma}}_{1},\xi^{\boldsymbol{\sigma}}_{2},\ldots,\xi^{\boldsymbol{\sigma}}_{n}\right)$ is its reordered version. Lastly, $\mathbf{Y}^{\boldsymbol{\sigma}}$ is the adjacency matrix reordered by $\boldsymbol{\sigma}$ for the columns and rows of $\mathbf{Y}$ simultaneously, such that $Y_{pq}^{\boldsymbol{\sigma}}=Y_{\sigma_{p}\sigma_{q}}$.
@@
The central component of the SBM is the distribution assumption about each dyad of $\mathcal{G}$, essentially each element of $\mathbf{Y}$ or equivalently $\mathbf{Y}^{\boldsymbol{\sigma}}$. As $\mathbf{Y}^{\boldsymbol{\sigma}}$ is upper triangular for a $\boldsymbol{\sigma}$ that is topological, we require that, for all dyads $(p,q)$ where $1\leq p<q\leq n$,
@@
$Y^{\boldsymbol{\sigma}}_{qp}=0,\quad\quad Y^{\boldsymbol{\sigma}}_{pq}|\mathbf{C},\mathbf{Z}_{n},\boldsymbol{\sigma},\boldsymbol{\xi}$ $\sim\text{Poisson}\left(\xi^{\boldsymbol{\sigma}}_{p}\xi^{\boldsymbol{\sigma}}_{q}C_{Z^{\boldsymbol{\sigma}}_{p}Z^{\boldsymbol{\sigma}}_{q}}\right),$ (2)
where $\mathbf{C}=(C_{ij})\in\mathbb{R}_{+}^{K_{n}\times K_{n}}$ is the block matrix.
As $\mathbf{Z}_{n}$ is unknown prior to fitting the model, it will be treated as a vector of latent variables and assigned a prior. Specifically, we assume that $\mathbf{Z}_{n}$ are the first of $n$ elements of an exchangeable sequence driven by a PY process prior. The parametrization of this prior will be detailed in Section 2.2. Given this prior choice, the vector $\mathbf{Z}_{n}$ will display $K_{n}$ distinct values $(Z_{1}^{*},\ldots,Z_{K_{n}}^{*})$, which appear in $\mathbf{Z}_{n}$ with frequencies $\mathbf{N}=(N_{1},\ldots,N_{K_{n}})$. The $i$-th element $N_{i}=\sum_{p=1}^{n}\mathbb{I}(Z_{p}=Z_{i}^{*})$ is the number of vertices in group $i$, displaying label $Z_{i}^{*}$. As previously mentioned, we write $\mathbb{I}(Z_{p}=i)$, in place of $\mathbb{I}(Z_{p}=Z_{i}^{*})$. Also, we derive the $K_{n}\times K_{n}$ edge matrix between groups from $\mathbf{Z}_{n}$ and $\mathbf{Y}$, denoted by $\mathbf{E}$, where $E_{ij}=\sum_{p=1}^{n}\sum_{q=1}^{n}Y_{pq}\mathbb{I}(Z_{p}=i,Z_{q}=j)$.
Next, we define a $K_{n}\times{}K_{n}$ matrix $\mathbf{M}$, where $M_{ij}$ is the number of dyads $(p,q)$ between groups $i$ and $j$ such that vertex $p$ is topologically in front of vertex $q$. Mathematically,
$M_{ij}=\sum_{p=1}^{n-1}\sum_{q=p+1}^{n}\xi^{\boldsymbol{\sigma}}_{p}\xi^{\boldsymbol{\sigma}}_{q}\mathbb{I}\left(Z_{\sigma_{p}}=i,Z_{\sigma_{q}}=j\right)=\sum_{p=1}^{n-1}\sum_{q=p+1}^{n}\xi^{\boldsymbol{\sigma}}_{p}\xi^{\boldsymbol{\sigma}}_{q}\mathbb{I}\left(Z^{\boldsymbol{\sigma}}_{p}=i,Z^{\boldsymbol{\sigma}}_{q}=j\right).$

### v3

@@ Abstract
Random graphs have been widely used in statistics, for example in network analysis and graphical models. In some applications, the data may contain an inherent hierarchical ordering among its vertices, which prevents directed edges between pairs of vertices that do not respect this order. For example, in bibliometrics, older papers cannot cite newer ones. In such situations, the resulting graph forms a Directed Acyclic Graph. In this article, we extend the Stochastic Block Model (SBM) to account for the presence of such ordering in the data, ignoring which can lead to biased estimates of the number of blocks. The proposed approach includes in the model likelihood a topological ordering, which is treated as an unknown parameter and endowed with a prior distribution. We describe how to formalise the model and perform posterior inference for a Bayesian nonparametric version of the SBM in which both the hierarchical ordering and the number of latent blocks are learnt from the data. Finally, an illustration with real-world datasets from bibliometrics is presented. Additional supplementary materials are available online.
@@ 1 Introduction
Random graphs have been widely used in statistics and computer science over the last years to model network, interaction and bibliometric data, just to mention a few among many fields of applications. For example, in social networks, friendships or message exchanges among users can be modelled as random graphs, having users as vertices, and edges between two users if they are friends or have exchanged messages. As another example, citation datasets in bibliometrics analysis can be modelled as random graphs, having articles as vertices, and directed edges from citing articles to cited articles.
@@
In some applications, the dataset and the corresponding graph can contain an inherent hierarchical order among its vertices. This means that directed edges can be present only from vertices that appear earlier in this order to ones that appear later. As an example, when modelling citation networks among journal articles, e.g. Ji and Jin (2016), articles can cite only articles that were published earlier. As another example, in a network having the employees of a big company as vertices and directed edges from the supervisor to the supervisee, a directed path from one vertex to another can be present only if the former is higher than the latter one in the company hierarchy. Similarly, the same structure appears in a dataset having directed edges from PhD supervisors to their students. In all these applications, the observed graph is a Directed Acyclic Graph (DAG), i.e. it cannot display any directed cycles (any directed path having the same initial and final vertex). Moreover, it respects a latent topological ordering among its vertices, i.e. a linear ordering $\prec$ of the vertex set $\mathcal{V}$ such that every pair $(p,q)\in\mathcal{V}\times\mathcal{V}$ can belong to the edge set $\mathcal{E}$ only if $p\prec q$. In general, the topological ordering is not unique, i.e. for a DAG there can be multiple orderings that satisfy the definition above.
@@
When analysing a given network dataset that is a DAG, if the latent topological ordering is not properly accounted for, the statistical models can assign positive probability to ‘impossible’ edges, i.e. directed edges in which the first vertex appears later than the second one in the topological ordering. If this is the case for many edges, the resulting estimates from the model can deteriorate dramatically. The same phenomenon arises other statistical contexts, for example in presence of structural zeros when modelling tabular data (Manrique-Vallier and Reiter, 2014), in which assigning, whatever small, positive probability to ‘impossible’ cells can deteriorate estimates and bias results.
@@
In this work, we propose an approach to extending the popular Stochastic Block Model (SBM) (Holland et al., 1983) for random graphs in general to graphs that are specifically DAGs. The SBM, being among the most popular random graph models for vertex clustering and community detection, was initially introduced by Holland et al. (1983) and Wang and Wong (1987), then formalised as latent models by Snijders and Nowicki (1997) and Nowicki and Snijders (2001), and has further gained popularity with the incorporation of degree correction by Karrer and Newman (2011). In online Appendices A.1 and A.2, we provide a brief review of this model, its generalisations and approaches to posterior inference. For recent comprehensive reviews, the reader is also referred to Abbe (2018) and Lee and Wilkinson (2019), which focus on theoretical results and modelling approaches, respectively. When extending the SBM to model a given DAG, the underlying topological ordering is unknown and not necessarily unique. Furthermore, ignoring the ordering can lead to dramatic underestimates or overestimates of the number of blocks, as the simulations in Section 3.1 demonstrate. Therefore, it will be included in the likelihood of the model as a parameter, endowed with a prior and inferred a posteriori, together with the community memberships and other parameters of the model.
@@
The background literature regarding the use of the SBM for DAGs is quite limited and mainly consists of an unpublished technical report by Lee and Wilkinson (2018), which only presents a preliminary data analysis of a DAG using a SBM. Not only was degree correction not incorporated, but there was also no formal consideration of the issue of the number of blocks. Instead, the model was merely fitted with different numbers of blocks separately with no model selection, thus not justifying the interpretations of their results.
@@
The issue of the number of communities or blocks is common with SBMs and clustering in general, and has been reviewed in Table 1 of Lee and Wilkinson (2019). One common approach to this issue is the use of Bayesian nonparametrics, which assigns a prior for the allocation vector and allows the posterior sampler to infer the number of blocks from the data. Kemp et al. (2006) and Schmidt and Mørup (2013) initially considered the infinite relational model, which is a variant of the plain SBM, using a Dirichlet process prior, while Geng et al. (2019) extends this model by the use of a finite regime Pitman-Yor (PY) process (Pitman and Yor, 1997), which is reviewed in the online Appendix A.3. A further generalisation is in Legramanti et al. (2022), in which the Gibbs-type prior formulation is used instead of the PY prior, for an application to criminal networks.
@@
In the proposed model, we also use a Bayesian nonparametric version of the SBM, but by adopting the PY process in a different way. This process allows two possible regimes: one in which the total number of blocks is finite (and unknown) and another one in which it increases with the sample size. While other works such as Geng et al. (2019) have used Bayesian nonparametric formulation for the block membership, they usually focused only on the first of the two regimes. As these two regimes imply different asymptotic behaviours of the number of blocks, the model might be misspecified if only one regime is assumed before fitting to the data. To circumvent this, the choice of regime becomes another part of the proposed model, with a model selection step based on Gibbs variable selection (Carlin and Chib, 1995) embedded in the Markov chain Monte Carlo (MCMC) sampler.
@@
The proposed approach is illustrated using both simulated data and real datasets from bibliometrics. In the simulation study, where the proposed model is compared to a SBM for directed graphs, which does not account for the latent order in the likelihood, the latter brings about the aforementioned issue of assigning positive probability to many ‘impossible’ edges, as well as substantial biases in the estimates of the overall number of blocks. In the applications to two datasets of citation networks, it is shown that the model can infer the community structure and latent hierarchical order among vertices.
@@
The rest of this article is organised as follows. The SBM for DAGs is presented in Section 2. Section 3 presents both simulated (Section 3.1) and real data (Sections 3.2 and 3.3) illustrations. A final discussion in Section 4 concludes the article. The online appendices contain reviews of the SBM and PY process, a detailed description of the MCMC sampler to perform posterior inference for the SBM for DAG, additional information on the likelihood derivations, and plots from the real-data illustrations.
@@ 2 Model
To introduce the model and notation used throughout the rest of the paper, let us consider a directed network to be represented as a directed multi-graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the vertex set and $\mathcal{E}$ is the edge set. The size of $\mathcal{V}$, denoted by $n:=|\mathcal{V}|$, is the number of vertices of $\mathcal{G}$. The $n\times{}n$ adjacency matrix of the graph is denoted by $\mathbf{Y}:=(Y_{pq})_{1\leq p,q\leq n}$. If there are $y$ directed edges from vertex $p$ to vertex $q$, i.e. $y$ copies of $(p,q)$ in $\mathcal{E}$, then $Y_{pq}=y$, otherwise $Y_{pq}=0$. We also assume no self-loops, i.e. $Y_{pp}=0$ for all $p=1,2,\ldots,n$. Next, we define $\mathbf{Z}_{n}:=(Z_{1},\ldots,Z_{n})$ to be the allocation vector of length $n$, where $Z_{p}$ is a label associated to vertex $p$. Two vertices, $p$ and $q$, belong to the same group if and only if $Z_{p}=Z_{q}$. Essentially, $\mathbf{Z}_{n}$ represents the group memberships of the vertices. We assume there are $K_{n}>1$ unique labels displayed in $\mathbf{Z}_{n}$, denoted $(Z_{1}^{*},\ldots,Z_{K_{n}}^{*})$ hence $K_{n}$ groups of vertices. When whether $Z_{i}^{*}$ is the label associated with vertex $p$ is of concern, for notational convenience, hereafter we write $\mathbb{I}(Z_{p}=i)$ in place of $\mathbb{I}(Z_{p}=Z_{i}^{*})$, where $\mathbb{I}(A)$ denotes the indicator function of the event $A$, i.e. $\mathbb{I}(A)=1$ if $A$ is true, 0 otherwise. We also define $\mathbf{C}:=(C_{ij})_{1\leq{}i,j\leq{}K_{n}}$ to be the block matrix, where $C_{ij}$ represents the general connectivity between blocks $i$ and $j$. Similar to above, given $\mathbf{C}$ and $Z_{p}=Z_{k}^{*}$, $C_{Z_{p}j}$ will denote element $C_{kj}$.
@@
The main idea of the classic SBM is that the number of edges $Y_{pq}$ from vertex $p$ to vertex $q$ is independent of that of any other dyad, conditional on their group memberships $Z_{p}$ and $Z_{q}$ (Holland et al., 1983, Nowicki and Snijders, 2001). Moreover, we will consider the degree-corrected version of the SBM (Karrer and Newman, 2011), which takes into account the degree heterogeneity of the vertices within the same group, by introducing the vector $\boldsymbol{\xi}=(\xi_{1},\ldots,\xi_{n})$ of vertex-specific parameters, where $\xi_{p}>0$ is the degree correction factor for vertex $p$. Namely, the number of edges between vertex $p$ and $q$ is modelled as
$Y_{pq}|\mathbf{C},\mathbf{Z}_{n},\boldsymbol{\xi}~\overset{\text{ind}}{\sim}~\text{Poisson}(\xi_{p}\xi_{q}C_{Z_{p}Z_{q}}).$ (1)
@@
To introduce the proposed model, we restrict $\mathcal{G}$ to be a DAG for the rest of this section. To utilise a unique feature of DAGs, as discussed in Section 1, we define $\boldsymbol{\sigma}:=(\sigma_{1},\ldots,\sigma_{n})$ as the random vector that represents the ordering of $\mathcal{G}$, with the collection of all permutations of $\{1,2,\ldots,n\}$ as the sample space. A value of $\boldsymbol{\sigma}$ is deemed topological with respect to $\mathcal{G}$ if it satisfies the definition of topological ordering, i.e. if vertex $p$ precedes vertex $q$ in the ordering, then there cannot be edges from $q$ to $p$. We define several quantities implied from $\boldsymbol{\sigma}$. First, $\boldsymbol{\phi}:=(\phi_{1},\ldots,\phi_{n})$ is the “inverse” of $\boldsymbol{\sigma}$, which means that if vertex $p$ is the $r$-th vertex in the topological ordering, we have $\sigma_{r}=p$ and $\phi_{p}=r$. Essentially, $\boldsymbol{\phi}$ contains the position of each vertex in $\boldsymbol{\sigma}$. Second, we define $\mathbf{Z}_{n}^{\boldsymbol{\sigma}}:=\left(Z^{\boldsymbol{\sigma}}_{1},Z^{\boldsymbol{\sigma}}_{2},\ldots,Z^{\boldsymbol{\sigma}}_{n}\right)$ as the reordered allocation vector, where $Z^{\boldsymbol{\sigma}}_{p}=Z_{\sigma_{p}}$ and $Z_{\sigma_{p}}$ comes from $\mathbf{Z}_{n}$. In the same way, we define $\boldsymbol{\xi}^{\boldsymbol{\sigma}}:=\left(\xi^{\boldsymbol{\sigma}}_{1},\xi^{\boldsymbol{\sigma}}_{2},\ldots,\xi^{\boldsymbol{\sigma}}_{n}\right)$ as the reordered version of $\boldsymbol{\xi}$. Lastly, $\mathbf{Y}^{\boldsymbol{\sigma}}$ is the adjacency matrix reordered by $\boldsymbol{\sigma}$ for the columns and rows of $\mathbf{Y}$ simultaneously, such that $Y_{pq}^{\boldsymbol{\sigma}}=Y_{\sigma_{p}\sigma_{q}}$.
@@
The central component of the SBM is the distribution assumption about each dyad of $\mathcal{G}$, essentially each element of $\mathbf{Y}$. Due to the acyclic natrue of $\mathcal{G}$, the combinations of the possible values of $(Y_{pq},Y_{qp})$ are restricted to be $(0,0)$, $(0,y)$ or $(y,0)$, for some positive integer $y\in\mathbb{N}$. Equivalently, as $\mathbf{Y}^{\boldsymbol{\sigma}}$ is upper triangular for a $\boldsymbol{\sigma}$ that is topological, we require that, for all dyads $(p,q)$ where $1\leq p<q\leq n$,
@@
$Y^{\boldsymbol{\sigma}}_{qp}=0,\quad\quad Y^{\boldsymbol{\sigma}}_{pq}|\mathbf{C},\mathbf{Z}_{n},\boldsymbol{\sigma},\boldsymbol{\xi}$ $\overset{\text{ind}}{\sim}\text{Poisson}\left(\xi^{\boldsymbol{\sigma}}_{p}\xi^{\boldsymbol{\sigma}}_{q}C_{Z^{\boldsymbol{\sigma}}_{p}Z^{\boldsymbol{\sigma}}_{q}}\right).$ (2)